OpenStudy (anonymous):

Help please, Algebra 1- Solving Rational Expressions 2c 4 --- - 2 = ----- c-4 c + 5

OpenStudy (anonymous):

you have to put everything with the same denominator first. Then i'll help you with the remaining equation

OpenStudy (anonymous):

Try multiplying both sides by \((c-4)(c+5)\) and simplify a bit.

OpenStudy (anonymous):

\[\frac{ 2c }{ c-4 }-2=\frac{ 4 }{c+5 }\] \[\frac{ 2c-2\left( c-4 \right) }{c-4 }=\frac{ 4 }{ c+5 }\] \[\frac{ 2c-2c+8 }{ c-4 }=\frac{ 4 }{ c+5 }\] \[or \frac{ 8 }{ c-4 }=\frac{ 4 }{ c+5 }\] \[8\left( c+5 \right)=4\left( c-4 \right)\] or 8c+40=4c-16 or 8c-4c=-16-40 or 4c=-56 or c=-56/4=-14